OpenStudy (anonymous):

Differentiate the function f(x) = sin x ln(5x)

OpenStudy (anonymous):

poduct plus chain rule

OpenStudy (anonymous):

\[(fg)'=f'g+g'f\] use \[f(x)=\sin9x), f'(x)=\cos(x), g(x)=\ln(5x), g'(x)=\frac{1}{x}\]

OpenStudy (anonymous):

cosxln5x + sinx/x

OpenStudy (anonymous):

typo there \[f(x)=\sin(x)\]

OpenStudy (anonymous):

the function is f(x) = sin(x) * ln(5x)

OpenStudy (anonymous):

the derivative of ln(5x) is 5/x or 1/5x?

OpenStudy (anonymous):

None. Use chain rule = (1/5x)*5= 1/x